In a sonometer experiment resonance is obtained when the experiment wire has a length of 21 CM between the bridges and the vibrations are excited by A tuning fork of frequency 256 Hertz. If A tuning fork of frequency 384 hertz is used what should be the length of the experimental wire to get the resonance.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Student
 
Here, l1 =21 cm, f1 = 256 Hz
l2 = ? , f2 = 384 Hz
As 
l1 * f1 = l2 * f2
l2 = 256 * 21/384 = 14 cm
 
Regards
Arun (AskIITians forum expert)